The coefficient of static friction between the tires of a car and a dry road is 0.62. The mass of the car is 1500 kg. What maximum braking force is obtainable (a) on a level road and (b) on an 8.6° downgrade?

Grade:12th pass

1 Answers

Vishal Mehta
7 years ago
u=.62, W=1500 (a) ∆=0° so F=uR= uWcos∆=.62×1500×1=930 N(b) ∆=8.6° so F= uR+Wsin∆=uWcos∆+Wsin∆= W(ucos8.6+sin8.6)=1500(.62(-.67)+.73)=1500×1.14=1718.1N